Suppose X is a random variable that has density function f(x) = (1/2)e^−|x| for −∞ < x < ∞. Find:
(a) (2 pts) P(X < 10).
(b) (4 pts) The c.d.f. of X2.
(c) (4 pts) V ar(X)
Answer #1

exo(u oo Soln 1 c 20 f(x) = exp(-1x1) 0 P(x <10) s fral du ja 4 0301-mo) din Sh £ (exp (2)] + Į (-exp(-ni] (1) + (1 1 - exp(-(ii) If no, then Fx (n = P(x <n) U I exp(-121) dn 2 1 s exp (2) du + & J exp(-aidh 2 e + = + [ exp hilo = 1 (1) + 4 1 - 1 expLet I s n exp(n) du expcns 10 [ n exp(n) 1 Iz = n exp(-) du exp (en)] [-sexpen) 1 +(*) (-1) co ㅋ E(X) 1 (2) = Simill as illa